periodcumsock:

A car manufacturer is reducing the number of incidents with the transmission by issuing a voluntary recall. During week 2 of the recall, the manufacturer fixed 192 cars. In week 4, manufacturer fixed 184 cars. Assume that the reduction in the number of cars each week is linear. Write an equation in function form to show the number of cars seen each week at the mechanic. f(x) = 4x + 200 f(x) = 2x + 192 f(x) = −4x + 200 f(x) = −2x + 192

BlazeNeedsHelp:

so what is it you need help with??

BlazeNeedsHelp:

like whats the exact part

periodcumsock:

i need help with the steps the whole thing confuses me

BlazeNeedsHelp:

okay what grade are you in and what math

Mercury:

let x = # of weeks, and y = # of cars, since the weeks is the independent variable and cars is dependent we can represent "During week 2 of the recall, the manufacturer fixed 192 cars. In week 4, manufacturer fixed 184 cars." as (2,192) and (4,184) now, we want a linear equation in y = mx + b form so we start by finding m. use the slope formula: \[m = \frac{ y_{2}-y_{1} }{ x_{2}-x_{1} }\] and plug in your points as (x1,y1) and (x2,y1). once you have m, take that m-value, and one of the original points, and plug into y = mx + b to solve for b. once you have m and b, re-write the original y = mx + b equation, plugging in m and b, and you should get your solution this is a lot, so please let me know if any step was confusing or if you just want to check your solution